what is the equation to find length and width if i am given just the perimeter? its harder because its in decimal form. i figured out the answer but i just need to know how to show my work
perimeter is 108.44in if the painting is 5.54in taller than it is wide, find dimensions of painting?
p= 108.44in
L= 5.54x in ?
w= ?
confused!

Answer by Maths68(1474)

5.54in taller than it is wide means we have to add 5.54 to its width then the length will be
l= w +5.54
w= ?
p=108.44
Perimeter = 2l + 2w
Plug in the value of l in above equation
108.44 = 2(w+5.54) + 2w
108.44 = 2w+11.08 + 2w
108.44 = 4w+11.08
108.44-11.08=4w
97.36=4w
Divide by 4 both sides of the equation
97.36/4=4w/4
24.34=w
w=24.34
l= 24.34 +5.54 [Plug in the value of w]
l=29.88
Dimensions of painting are
Length = 29.88 in
Width =24.34
Perimeter = 108.44